# Distributed Lloyd-Based Algorithm for Uncertainty-Aware Multi-Robot Under-Canopy Flocking

### v2

@@ Abstract
In this letter, we present a distributed algorithm for flocking in complex environments that operates at constant altitude, without explicit communication, no a priori information about the environment, and by using only on-board sensing and computation capabilities. We provide sufficient conditions to guarantee that each robot reaches its goal region in a finite time, avoiding collisions with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity constraint). The proposed approach allows to operate in crowded scenarios and to deal with tracking errors and on-board sensing errors, without violating safety and proximity constraints. The algorithm was verified through simulations with varying number of UAVs and also through numerous real-world experiments in a dense forest involving up to four UAVs.
@@
Video—https://mrs.fel.cvut.cz/rbl-wild-2024
Github—https://github.com/ctu-mrs/rbl_controller
@@ I Introduction
Over the past few decades, numerous applications have emerged for multi-robot systems, exhibiting their undisputed benefits in diverse fields. These applications span domains such as logistics, surveillance, environment monitoring, disaster response, search and rescue, and agriculture among others, demonstrating their versatility. For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization. The definition of flocking is provided by the seminal work of Reynolds [1], which defines it as a non-colliding aggregate motion. Depending on the environments where robots operate, the reliability or availability of explicit communication between robots may be compromised.
In this work, we propose a Lloyd-based algorithm that relies only on on-board sensors. We aim to synthesize a distributed algorithm, where each robot safely reaches the desired position, while maintaining proximity between robots in the team. We provide sufficient conditions to guarantee safety, proximity maintenance and convergence towards the goal in cluttered environments by accounting for bounded tracking errors and uncertainties in the estimation of neighbors and obstacles’ positions. Our algorithm takes inspiration from the nature; just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication. We validate our approach through simulations and experiments in the forest (see Figure 1).
@@ I-A Related works
To position our algorithm within the state of the art, we consider the following features. Distributed control: each robot makes decisions on the basis of local information, which is obtained by relying only on on-board sensors. Asynchronous control: each robot makes decisions independently, in asynchronous fashion. Without explicit communication: the robots do not exchange time-dependent information such as future trajectories, current positions or velocities. Moreover, they do not rely on a common reference frame. Safety: collision and obstacle avoidance is guaranteed at any time. Convergence: the sufficient conditions for the robots to converge to their goal locations are provided. Proximity maintenance: the distance between robots is guaranteed to be below the set threshold on maximum distance. Uncertainties: the system is provably safe and respects proximity constraints even in the presence of bounded tracking errors and/or bounded measurement uncertainties. Cluttered and highly dense environments: the algorithm can operate in cluttered environments such as a forest. In particular, in absence of uncertainties and convergence guarantees, the robots can operate in environments where the minimum distance between two obstacles is large enough to allow the single robot to pass. To ensure convergence and account for uncertainties, a sufficient condition is that the obstacles are spaced far enough to allow passage of twice the robot’s size. On-board localization: the algorithm is verified in unknown environments relying only on on-board localization. Experiments: the algorithm is tested on real robotic platforms in unstructured environments. Open source code: the algorithm is open source.
@@
Convergence $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
@@
In Table I, we report an overview of the most closely related algorithms in the literature. We mark with $\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially. As shown in Table I, our algorithm is unmatched in fulfilling all the specified features of interest. In fact, despite [5, 4] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network, the absence of both convergence analysis and a rigorous account of tracking errors and uncertainties. Similarly, in [6] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations. On the other hand, [7, 8], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety, proximity maintenance, convergence guarantees analysis, and robustness to tracking errors and uncertainties. In [10], the authors present a distributed solution, which in theory does not rely on explicit communication and accounts for uncertainties. However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions). With respect to [3, 27], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraints, i.e., two robots are considered in the proximity if they are close enough. This allows to increase the robots mobility and also to not rely on explicit communication. In Table  I, we report additional relevant state-of-the-art algorithms. Notably, all of them share limitations related to the use of an explicit communication network and/or to not rigorously account for safety, convergence, proximity maintenance and the uncertainties. Conversely, in our work, we provide a method which implements all the fore-mentioned features as well as a comprehensive rule for the selection of the algorithm’s parameters based on measurable quantities (e.g., tracking error, or covariance of state estimates), which is often an unclear or empirical process in the state-of-the-art algorithms.
@@ I-B Contribution and paper organization
Our contribution is threefold. First, we propose a novel algorithm for flocking control in dense cluttered environments. The key novelties lie in our use of an adaptive control strategy to compensate for tracking errors and measurement uncertainties, combined with Convex Weighted Voronoi Diagrams (CWVD) to enhance the performance in a dense forest scenarios. Second, we provide sufficient conditions for safety, proximity maintenance, and convergence in cluttered environments and in the presence of flocking constraints. Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar and UVDAR system [28], without relying neither on an explicit communication network nor GNSS signal.
@@ II Problem description
Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain connected the entire mission. We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma_{ij}=\Gamma_{ji},\,\,\,\forall i\in{1,\dots,N},\forall j\in\bar{\mathcal{N}}_{i}$. Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
@@ III Proposed Approach
where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [13], which accounts only for neighboring robots’ positions
@@
While $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles positions
@@
where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$, is a parameter, which regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
Figure 2: Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$. Where the green circle is the i–th robot and the cyan circles represents the neighboring robots.
In Figure 2, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
@@ III-A Reshaping the cell geometry $\mathcal{V}_{i}$
By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
@@
The proof comes from the definition of convex set. Due to the fact that the set $\mathcal{A}_{i}$ is always convex, and by definition $p_{i},c_{\mathcal{A}_{i}}\in\mathcal{A}_{i}$, it follows that since the straight path from the robot position to the centroid position is a safe path, then applying $\dot{p}_{i}(\mathcal{A}_{i})$ in (4) preserves safety. ∎
We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve swarming behavior. By imposing our final cell geometry $\mathcal{F}_{i}=\mathcal{A}_{i}\cap\mathcal{M}_{i}$, where
@@
we obtain proximity maintenance while preserving safety. In Figure 3, we depict an example of cell geometry $\mathcal{F}_{i}$.
Figure 3: Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$. Dashed lines indicate the proximity constraints. Obstacles are represented as grey circles, while the robots are depicted in cyan.
Theorem 2 (Proximity maintenance).
@@ III-B The weighting function $\varphi_{i}(q)$
Although ensuring safety and proximity maintenance only involves properly shaping the cell geometry, achieving convergence and enhancing navigation performance largely hinges on how the function $\varphi_{i}(q)$ is defined. We adopt a Laplacian-shaped function centered at $\bar{p}_{i}$, which represents a point in the mission space that ultimately needs to converge to the goal location. We design the $\varphi_{i}(q)$ function as follows:
@@
where the spreading factor $\beta_{i}\in\mathbb{R}$ is time-variant, it follows the dynamics
$\dot{\beta}_{i}(\mathcal{A}_{i})=\begin{cases}-\beta_{i}\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{1}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{2}\\ -(\beta_{i}-\beta^{D}_{i})&\text{otherwise,}\end{cases}$ (11)
@@
The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$.
Proposition 1 (Convergence in cluttered environment and with flocking constraints).
Let us assume we have an unbounded space cluttered with circular obstacles of radius $\delta^{o}<\Gamma_{ij}/2-2\delta$, where $\delta=\delta_{i},\forall i=1,\dots,N$, and at a minimum distance of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o}$ one to the other. The initial and goal positions are adequately apart from each other, i.e., $\Gamma_{ij}>\|p_{i}(0)-p_{j}(0)\|>\Delta_{ij}$, $\Gamma_{ij}\geq\|e_{i}-e_{j}\|>\Delta_{ij}$, $\forall i,j=1\dots N$ with $i\neq j$. By imposing the control law $\dot{p}_{i}(\mathcal{F}_{i})$ in (4), (10), (11), and (12) to compute the centroid position in (5), by properly selecting $d_{1},d_{2},d_{3},d_{4}$, and $\beta^{D}_{i}$, we have that $\lim_{t\rightarrow\infty}\|p_{i}(t)-e_{i}\|<r_{s,i},\forall i=1,\dots,N.$
Figure 4: The figure shows two possible deadlock conditions: case 1. is due to the small inter-space between obstacles. case 2. is due to the too short $\Gamma_{ij}$ value (or too big $\delta^{o}$ value). We depict in cyan the robots and in grey the static obstacles. The crosses represent the final goal positions while the dashed lines represent the proximity constraint topology.
Proof.
The proof provided in [2] does not consider presence of static obstacles nor proximity constraints. By including both of them, we recognize, beside the trivial initial and final conditions requirements, the addition of two potential deadlock scenarios, which are depicted in Figure 4.
Notice that the addition of static obstacles is equivalent to non-cooperative static robots. In order to not affect the proof of convergence obtained in [2], two requirements need to be considered. First, we require a minimum distance between the obstacles of $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$, in order to avoid the presence of narrow hallways that can generate deadlock issues (left hand side of Figure 4). This is a consequence of the results in [2]. In fact, static obstacles can be thought as robots that reached their goal and have a value of $\beta_{i}\rightarrow 0$. However, from [2], by design this is not possible if the robots are at a distance from their goal smaller than $d_{2}$. Since $d_{2}>2\delta$ and $\|e_{i}-e_{j}\|>\Delta_{ij}$, then the condition $\|o_{i}-o_{j}\|>4\delta+2\delta^{o},\,\forall i\neq j$ is equivalent. Second, we require that the maximum radius of the obstacles is $\delta^{o}<\Gamma_{ij}/2-2\delta$, in order to avoid strict proximity maintenance constraints (right hand side of Figure 4). By imposing these additional conditions, the guarantees of convergence in [2] holds. ∎

### v3

# arXiv:2504.18840v3[cs.RO] 27 Feb 2026

@@ Abstract
In this letter, we present a distributed algorithm for flocking in complex environments that operates at constant altitude, without explicit communication, no a priori information about the environment, and by using only on-board sensing and computation capabilities. We provide sufficient conditions to guarantee collision avoidance with obstacles and other robots without exceeding a desired maximum distance from a predefined set of neighbors (flocking or proximity maintenance constraint) during the mission. The proposed approach allows to operate in crowded scenarios and to explicitly deal with tracking errors and on-board sensing errors. The algorithm was verified through simulations with varying number of UAVs and also through numerous real-world experiments in a dense forest involving up to four UAVs.
@@
Video—https://mrs.fel.cvut.cz/rbl
Github—https://github.com/ctu-mrs/rbl
@@ I Introduction
Over the past few decades, numerous applications have emerged for multi-robot systems, exhibiting their undisputed benefits in diverse fields. These applications span domains such as logistics, surveillance, environment monitoring, disaster response, search and rescue, and agriculture among others, demonstrating their versatility. For many of these applications, having a swarm of robots able to flock together towards a desired location can be highly beneficial, due to redundancies and parallelization. Depending on the environments where robots operate, the reliability or availability of explicit communication between robots may be compromised.
In this work, we propose a Lloyd-based algorithm that relies solely on on-board sensing. We have synthesized a distributed algorithm in which each robot safely reaches its desired position while maintaining proximity to its teammates in cluttered environments. We choose Lloyd algorithm as a base for our method as it enables seamless multi-robot coordination without requiring robots to share unobservable or global information, and its safety guarantees are not dependent on finely tuned parameters.
Our algorithm takes inspiration from the nature; just like flock of birds or a school of fish, the robots do not need a leader and each robot makes its own decisions based on relative information, without relying on explicit communication. We validate our approach through simulations and experiments in the forest (see Figure 1).
@@ I-A Related works
To position our work within the state of the art, we consider the following features. Distributed control: each robot makes decisions on the basis of local information, which is obtained by relying only on on-board sensors. Asynchronous control: each robot makes decisions independently, in asynchronous fashion. Without explicit communication: the robots do not exchange time-dependent information such as future trajectories, current positions or velocities. Moreover, they do not rely on a common reference frame. Safety: collision and obstacle avoidance is guaranteed at any time. Proximity maintenance: the distance between robots is guaranteed to be below the set threshold on maximum distance. Uncertainties: the approach is robust to bounded tracking errors and bounded measurement uncertainties. Cluttered and highly dense environments: the algorithm can operate in cluttered environments such as a forest. On-board localization: the algorithm is verified in unknown environments relying only on on-board localization. Experiments: the algorithm is tested on real robotic platforms in unstructured environments. Open source code: the algorithm is open source.
@@
In Table I, we report an overview of the most closely related algorithms in the literature. We mark with $\times$” if the algorithm does not implement the feature, “✓” if it does, and finally, “o” if it does but only partially. As shown in Table I, our algorithm is unmatched in fulfilling all the specified features of interest. In fact, with respect to [9], we included the proximity maintenance constraints, the ability to deal with cluttered environments, tracking errors and uncertainties, without relying on explicit communication, nor an external localization system. Despite [30, 29] proposing valid solutions to the problem, their main limitations are the requirement of a trajectory broadcasting network and the absence of account of proximity maintenance, tracking errors and uncertainties. Similarly, in [31] the robots need to broadcast information such as poses, velocities, poses covariance, and global extrinsic transformations. On the other hand, [2, 17], do not rely on an explicit communication network, but they exhibit degraded performance in crowded scenarios, lack of safety guarantees, proximity maintenance, and robustness to tracking errors and uncertainties. In [32], the authors present a distributed solution, which in theory does not rely on explicit communication and accounts for uncertainties. However, in the experiments they rely on motion capture measurements adding limited noise errors, which is reasonable only in structured environment (i.e., laboratory conditions). With respect to [8, 5], where the authors introduced the flocking behavior in the Lloyd-based framework, we do not impose the line-of-sight constraints, i.e., two robots are considered in the proximity if they are close enough. This allows to increase the robots mobility and also to not rely on explicit communication. In Table I, we report additional relevant state-of-the-art algorithms. Notably, all of them share limitations related to the use of an explicit communication network and/or to not account for safety, proximity maintenance and the uncertainties. Conversely, in our work, we provide a method which implements all the fore-mentioned features as well as a comprehensive rule for the selection of the algorithm’s parameters based on measurable quantities (e.g., tracking error, or covariance of state estimates), which is often an unclear or empirical process in the state-of-the-art algorithms.
@@ I-B Contribution and paper organization
Our contribution is threefold. First, we propose a novel algorithm for flocking control in dense cluttered environments. The key novelties lie in our use of an adaptive control strategy to compensate for tracking errors and measurement uncertainties, combined with Convex Weighted Voronoi Diagrams (CWVD) to enhance the performance in a dense forest scenarios. Second, we provide sufficient conditions for safety and proximity maintenance in cluttered environments and in the presence of flocking constraints. Finally, we validate the algorithm in the forest, relying only on on-board sensors, i.e., Inertial Measurement Unit (IMU), 2-D Lidar, altitude sensor, and UVDAR system [26], without relying neither on an explicit communication network nor GNSS signals nor having a common reference frame.
@@ II Problem description
Let $N$ be the number of robots in the swarm, let $\mathcal{N}_{i}$ be the set of neighbors of the robot $i$, and $\bar{\mathcal{N}}_{i}\subseteq\mathcal{N}_{i}$ the set of neighbors that has to remain in the proximity of the robot $i$ for the entire mission. We want to synthesize a distributed control law, such that the swarm maintains the desired proximity constraints, which is characterized by a set of maximum distances $\Gamma=\{\Gamma_{ij}\mid i\in{1,\dots,N},j\in\bar{\mathcal{N}}_{i}\}$. Each robot in the swarm has to reach its goal region avoiding collisions with other robots and obstacles.
@@ III Proposed Approach
where $\mathcal{V}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{N}\}$, $\mathcal{V}_{i}=\mathcal{V}_{i}^{p}\cap\mathcal{V}_{i}^{o}$ with $\mathcal{V}_{i}^{p}$ being the ii–th Convex Weighted Voronoi Diagram (CWVD) introduced in [6], which accounts only for neighboring robots’ positions, while $\mathcal{V}_{i}^{o}$ is the ii–th CWVD accounting only for neighboring obstacles’ positions:
@@
where $w_{i}=[w_{i}^{x},w_{i}^{y}]^{\top}=q-p_{i}$ and $\alpha_{ij}=\arctan\left(\frac{y_{j}-y_{i}}{x_{j}-x_{i}}\right)$, while $\epsilon_{p(o)}\in[1,2]$ regulates the cautiousness in the robot motion (i.e., $\epsilon_{p(o)}=1$ more aggressive behavior, $\epsilon_{p(o)}=2$ more conservative behavior, notice that the latter coincides with the classic Voronoi diagram).
Refer to caption Refer to caption
(a) (b)
Figure 2: (a) Representation of the Convex Weighted Voronoi Diagram (CWVD) $\mathcal{V}_{i}^{p}$ in (2) with $\epsilon_{p}=2$ and $\epsilon_{p}=1$. Where the green circle is the i–th robot and the cyan circles represent the neighboring robots. (b) Example of $\mathcal{F}_{i}$ and $\mathcal{A}_{i}$ cell geometry with $\epsilon_{p}=\epsilon_{o}=2$. Red solid lines indicate the proximity maintenance constraints. Obstacles are represented as grey circles, while the robots are depicted in cyan.
In Figure 2a, we depict an example of the CWVD (2) for the case where $\epsilon_{p}=2$ and $\epsilon_{p}=1$.
@@ III-A Reshaping the cell geometry $\mathcal{V}_{i}$
By imposing the control $\dot{p}_{i}(\mathcal{A}_{i})$ in (4), with $\mathcal{A}_{i}=\tilde{\mathcal{V}}_{i}^{p}\cap\tilde{\mathcal{V}}_{i}^{o}\cap\mathcal{S}_{i}$, collision avoidance is guaranteed at every instant of time.
@@
The proof comes from the definition of convex set. Due to the fact that the set $\mathcal{A}_{i}$ is always convex, and by definition $p_{i},c_{\mathcal{A}_{i}}\in\mathcal{A}_{i}$, it follows that since the straight path from the robot position to the centroid position is a safe path, then applying $\dot{p}_{i}(\mathcal{A}_{i})$ in (4) preserves safety. Due to the robots’ size, the condition that the straight path from the robot to its centroid position is a safe path can be violated only if the $i$-th robot is at a distance greater than $2\Delta_{ij}$ ($2\Delta^{o}_{ij}$) from the $j$-th robot ($j$-th obstacle), according to (6), (7). Nevertheless, to experience collision the distance between the $i$-th robot and the $j$-th robot ($j$-th obstacle) needs to be less than $2\Delta_{ij}$ ($2\Delta^{o}_{ij})$, i.e., equal to $\Delta_{ij}$ ($\Delta^{o}_{ij}$), hence safety is preserved. ∎
We further apply an additional reshaping of the cell geometry in order to maintain the desired proximity of the agents, i.e., to achieve flocking behavior. By imposing our final cell geometry $\mathcal{F}_{i}=\mathcal{A}_{i}\cap\mathcal{M}_{i}$, where
@@
we obtain proximity maintenance while preserving safety. In Figure 2b, we depict an example of cell geometry $\mathcal{F}_{i}$.
Theorem 2 (Proximity maintenance or flocking constraint).
@@ III-B The weighting function $\varphi_{i}(q)$
Although ensuring safety and proximity maintenance only involves properly shaping the cell geometry, convergence to the goal and enhancing navigation performance largely hinges on how the function $\varphi_{i}(q)$ is defined. As originally proposed in [9], we adopt a Laplacian-shaped function centered at $\bar{p}_{i}$, which represents a point in the mission space that ultimately needs to converge to the goal location. We design the $\varphi_{i}(q)$ function as follows:
@@
where the spreading factor $\beta_{i}\in\mathbb{R}$ is time-variant. It follows the dynamics
$\dot{\beta}_{i}(\mathcal{A}_{i})=\begin{cases}-k_{\beta}\beta_{i}\,\,\,&\text{if}\,\,\|c_{\mathcal{A}_{i}}-p_{i}\|<d_{1}\,\land\\ &\|c_{\mathcal{A}_{i}}-c_{\mathcal{S}_{i}}\|>d_{2},\\ -k_{\beta}(\beta_{i}-\beta^{D}_{i})&\text{otherwise,}\end{cases}$ (11)
@@
The parameters $\beta_{i}^{D},d_{1},d_{2},d_{3},d_{4},k_{\beta},k_{e}$ are positive scalars, $R(\theta)$ is the rotation matrix, $e_{i}$ is the goal position for the ii–th robot, $\varepsilon$ is a small number, which allows to avoid rotations equal to $\pi/2$, while $\bar{{c}}_{\mathcal{A}_{i}}$ is the centroid position computed over the cell $\mathcal{A}_{i}$ with $\bar{p}_{i}\equiv e_{i}$. Equations (11) and (12) specify two adaptive rules. Intuitively, (11) regulates the greediness of the robot $i$ in reaching its goal region, while (12) introduces a right (or left) hand-side rule to enforce a common behavioral convention across the robots and hence enhance inter-robot coordination.